OpenStudy (anonymous):

A company that sells computer keyboards has a revenue of 325x - 3x^2 dollars and a profit of 3x^2 - 255x - 250 for every keyboard x. If Profit = Revenue - Cost, find the company's cost equation

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Profit = Revenue - Cost (3x^2 - 255x - 250) = (325x - 3x^2 ) - C Now solve for C

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

umm idk

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

I'm listening to really disturbing music that's perverted and hardcore, I'll get to you when I can ok?

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

A company that sells computer keyboards has a revenue of 325x - 3x^2 dollars and a profit of 3x^2 - 255x - 250 for every keyboard x. If Profit = Revenue - Cost, find the company's cost equation. \[3x ^{2}-255x-250=325x-3x ^{2}-C\]+3x^2 +255x +255x+3x^2 \[6x ^{2} - 250 = 580x - C\]

OpenStudy (anonymous):

6x^2 - 580x - 250 = -C C = -6x^2 +580x + 250

OpenStudy (anonymous):

C should equal C = -6x^2 +580x + 250

OpenStudy (anonymous):

Are those any of your choices or is it not multiple choice?

OpenStudy (anonymous):

@Cosmichaotic